2 2 2
35in x (2.54 cm) x (1m)
-------- -------- 2 ------ 2
1 (1 in) (100 cm)

Answer by Alan3354(69443)

What does that mean?
-----
35 in = 35*2.54 cm = 35*2.54/100 meters
What are the 2's about?